Please assist,
4x-y=16
x+6y=29
Thank you!

Answer by gwendolyn(128)

The second equation states:
x+6y=29
We can manipulate this equation to find x in terms of y by subtracting 6y from both sides of the equation:
x+6y=29
x+6y-6y=29-6y
x=29-6y
We can substitute this value for x into the first equation in the problem.
4x-y=16
4(29-6y)-y=16
Next, we can distribute the 4 over the first set of parentheses.
116-24y-y=16
116-25y=16
After that, we can subtract 116 from both sides of the equation.
-25y=-100
Finally, we divide both sides of the equation by negative 25 to isolate our variable.
y=4
All that we need to do now is find the value of x:
x=29-6y
x=29-24
x=5
So, y=4 and x=5